OpenStudy (anonymous):

Which of the following is a rational number? square root 15 square root 16, square root 21, square root 22

OpenStudy (anonymous):

@some.random.cool.kid

OpenStudy (some.random.cool.kid):

hold on almost done.

OpenStudy (cwrw238):

is any one of the numbers 15, 16 21 or 22 a perfect square?

OpenStudy (anonymous):

Yes

OpenStudy (some.random.cool.kid):

ok so rational means that it can be divided by 1 if it can be then its rational so thats how you can solve this

OpenStudy (some.random.cool.kid):

also it would have to be a whole number

OpenStudy (anonymous):

So would 15 be the answer?

OpenStudy (cwrw238):

no

OpenStudy (some.random.cool.kid):

yeah no it would be ... something with uh how to put this

OpenStudy (cwrw238):

start from 1 multiply each number by 4...

OpenStudy (cwrw238):

1* 4 = 4 . 2 * 4 = ?

OpenStudy (anonymous):

I don't know how to do all that. But I do think the answer is 15.

OpenStudy (anonymous):

@cwrw238 great idea

OpenStudy (some.random.cool.kid):

well no its not 15

OpenStudy (anonymous):

8

OpenStudy (some.random.cool.kid):

your close though :} its higher

OpenStudy (anonymous):

16

OpenStudy (cwrw238):

thats 4*2 - carry on 4*3 = ?

OpenStudy (cwrw238):

right 16 is a perfect square , 4 times 4 = 16

OpenStudy (some.random.cool.kid):

yay

OpenStudy (some.random.cool.kid):

correct

OpenStudy (some.random.cool.kid):

:} great job

OpenStudy (cwrw238):

square root of 16 is 4 , 4 is rational

OpenStudy (anonymous):

THANKS!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Can you guys help me with a few more questions

OpenStudy (anonymous):

there you go! @timeflybyegirl @cwrw238 was trying to get you to calculate that so you can get the answer!

OpenStudy (some.random.cool.kid):

no prob bye bye its been fun wroking with you timefly but i kind of i have to finish my own math module exam so bye.

OpenStudy (anonymous):

Bye thanks for helping :)